Net capacitance of three identical capacitors connected in parallel is 12 microfarad. What will be
the net capacitance when two of them are connected in
(i) parallel
(ii) series?

Grade:12

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
In parallel combination,C=C_{1}+C_{2}+C_{3}
Also, since the capacirors are identical, So each is 4 microfarad.
and hence when two of them are connecetd in parallel, the total capacitance is 4+4=8 microfarad.
and when both are in series, the net capacitance is given by
1/C=1/C_{1}+1/C_{2}
so C= 2 microfarad.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty